# Byzantine-resilient Federated Learning With Adaptivity to Data Heterogeneity

### v3

@@ VI Conclusion
In this paper, we propose a novel FL aggregation strategy, RAGA, so as to be robust to Byzantine attacks and data heterogeneity. In our proposed RAGA, geometric median is leveraged for aggregation and the round number for local updating can be selected freely. Rigorous proofs disclose that our proposed RAGA can achieve convergence at rate $\mathcal{O}({1}/{T^{2/3-\delta}})$ with $\delta\in(0,2/3)$ for non-convex loss function and at linear rate for strongly-convex loss function, so long as the fraction of dataset from malicious users is less than half. Experimental results based on heterogeneous datasets show that RAGA performs steadily and keeps on an advantage on convergence performance over baseline methods under various intensity of Byzantine attacks.
Appendix A Proof of Theorem 1
@@
With Assumptions 1, 3, 4, 5, and $C_{\alpha}>0.5$, the term $\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$ can be upper bounded as
$\displaystyle\quad\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}\leqslant\Delta^{t}+\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (29)
Proof:
Please refer to Appendix C. ∎
Lemma 2
With Assumptions 3, 4, 5, and $C_{\alpha}>0.5$, the term $\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}$ can be bounded as
$\displaystyle 0\leqslant\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (30)
@@
$\displaystyle=\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}-\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{(\eta^{t})^{2}L-\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}$ (31c)
$\displaystyle=\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}-\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{(\eta^{t})^{2}L-\eta^{t}}{2}p^{t}\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}$ (31d)
$\displaystyle\leqslant-\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{\eta^{t}}{2}\Delta^{t}+\frac{(\eta^{t}+p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})\epsilon^{2}}{(2C_{\alpha}-1)^{2}}+\frac{4(C_{\alpha})^{2}(\eta^{t}\sigma^{2}+(p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})(G^{2}+\sigma^{2}))}{(2C_{\alpha}-1)^{2}}$ (31e)
@@
$\displaystyle\quad+\sum_{t=1}^{T}\frac{2(\eta^{t}+p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})\epsilon^{2}}{(2C_{\alpha}-1)^{2}\sum_{t^{\prime}=1}^{T}\eta^{t}}+\sum_{t=1}^{T}\frac{8(C_{\alpha})^{2}(\eta^{t}\sigma^{2}+(p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})(G^{2}+\sigma^{2}))}{(2C_{\alpha}-1)^{2}\sum_{t^{\prime}=1}^{T}\eta^{t}}$ (32)
@@
Appendix B Proof of Theorem 2
@@
$\displaystyle=\frac{L}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{t+1}-\bm{w}^{*}\right\rVert^{2}\right\}$ (33b)
For the term $\mathbb{E}\left\{\left\lVert\bm{w}^{t+1}-\bm{w}^{*}\right\rVert^{2}\right\}$, with $1\displaystyle 0<\lambda^{t}<$ 1, it leads to
@@
$\displaystyle=\mathbb{E}\Bigl{\{}\big{\lVert}\bm{w}^{t+1}-\bm{w}^{t}+\eta^{t}\nabla F(\bm{w}^{t})+\bm{w}^{t}-\eta^{t}\nabla F(\bm{w}^{t})-\bm{w}^{*}\big{\rVert}^{2}\Bigr{\}}$ (34a)
$\displaystyle\leqslant\frac{1}{\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t+1}-\bm{w}^{t}+\eta^{t}\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{1}{1-\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t}-\eta^{t}\nabla F(\bm{w}^{t})-\bm{w}^{*}\right\rVert^{2}\right\}$ (34b)
$\displaystyle=\frac{(\eta^{t})^{2}}{\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{1}{1-\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t}-\bm{w}^{*}\right\rVert^{2}\right\}-\frac{2\eta^{t}}{1-\lambda^{t}}\left\langle\bm{w}^{t}-\bm{w}^{*},\nabla F(\bm{w}^{t})-\nabla F(\bm{w}^{*})\right\rangle$
@@
$\displaystyle\leqslant\frac{(\eta^{t})^{2}}{\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{1-2\mu\eta^{t}+L^{2}(\eta^{t})^{2}}{1-\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t}-\bm{w}^{*}\right\rVert^{2}\right\}$ (34d)
@@
  • the inequality (34d) is bounded by Assumptions 1 and 2.
@@
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{w}^{T}-\bm{w}^{*}\right\rVert^{2}\right\}\leqslant\prod_{t=1}^{T-1}\gamma^{t}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}+\sum_{t=1}^{T-1}\frac{(\eta^{t})^{2}}{\lambda^{t}}\prod_{i=t}^{T-1}(\gamma^{i+1})^{q^{i+1}}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$ (35)
With the help of Lemma 1, and combining (33) and (35), there is
@@
$\displaystyle\leqslant\frac{L}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}\prod_{t=1}^{T-1}\gamma^{t}+\frac{L}{2}\sum_{t=1}^{T-1}\frac{(\eta^{t})^{2}}{\lambda^{t}}\Bigg{[}\Delta^{t}+\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}\Bigg{]}\cdot\prod_{i=t}^{T-1}(\gamma^{i+1})^{q^{i+1}}$ (36)
@@
Appendix C Proof of Lemma 1

### v4

@@ VI Conclusion
In this paper, we propose a novel FL aggregation strategy, RAGA, designed to be robust against Byzantine attacks and data heterogeneity. Our proposed RAGA leverages the geometric median for aggregation, allowing flexible selection of the round number for local updates. Rigorous proofs reveal that RAGA can achieve a convergence rate of $\mathcal{O}({1}/{T^{2/3-\delta}})$ for non-convex loss functions, with $\delta\in(0,2/3)$, and a linear rate for strongly-convex loss functions, provided that the fraction of the dataset from malicious users is less than half. Experimental results on heterogeneous datasets demonstrate that RAGA consistently outperforms baseline methods in convergence performance under varying intensities of Byzantine attacks.
Appendix A Assumptions And Theoretical Results
In this section, we provide assumptions and theoretical results of the RAGA under Byzantine attacks. Below, we first present the necessary assumptions.
A-A Assumption
For all honest users $m\in\mathcal{M}\setminus\mathcal{B}$, we have the following assumptions.
Assumption 7 (Lipschitz Continuity)
$\displaystyle f(\bm{w}_{1},\bm{s})-f(\bm{w}_{2},\bm{s})$ $\displaystyle\leqslant\left\langle\nabla f(\bm{w}_{2},\bm{s}),\bm{w}_{1}-\bm{w}_{2}\right\rangle$
$\displaystyle\qquad\qquad\quad+\frac{L}{2}\left\lVert\bm{w}_{1}-\bm{w}_{2}\right\rVert^{2},$ (30)
which is equivalent to the following inequality,
With (7) or (31), it can be derived that the global loss function $F(\bm{w})$ and local loss function $F_{m}(\bm{w})$ both have $L$-Lipschitz continuity.
Assumption 8 (Strongly-Convex)
Similarly, with (32), $F(\bm{w})$ and $F_{m}(\bm{w})$ can be derived to be both $\mu$ strongly-convex.
Assumption 9 (Local Unbiased Gradient)
Suppose a subset of dataset $\mathcal{S}_{m}$ is selected randomly, which is denote as $\xi_{m}$, define
the unbiased gradient [37, 38] implies that
$\mathbb{E}\left\{\nabla F_{m}(\bm{w};\xi_{m})\right\}=\nabla F_{m}(\bm{w}).$ (34)
Assumption 10 (Bounded Inner Variance)
For each honest user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the variance of stochastic local gradient $\nabla F_{m}(\bm{w};\xi_{m})$ is upper-bounded [37, 38] by
$\displaystyle\text{Var}(\nabla F_{m}(\bm{w};\xi_{m}))\leqslant\sigma^{2}.$ (35)
Assumption 11 (Data Heterogeneity)
For each honest user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the data heterogeneity can be defined as
$\displaystyle\theta_{m}=\left\lVert\nabla F_{m}(\bm{w})-\nabla F(\bm{w})\right\rVert.$ (36)
We also assume the data heterogeneity is bounded [37, 38], which implies
$\displaystyle\theta_{m}\leqslant\theta.$ (37)
Assumption 12 (Bounded Gradient)
For each honest user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the ideal local gradient $\nabla F_{m}(\bm{w})$ is upper-bounded [37, 38] by
$\displaystyle\left\lVert\nabla F_{m}(\bm{w})\right\rVert\leqslant G.$ (38)
A-B Theoretical Results
In this subsection, we present the analytical results of convergence on our proposed RAGA. All proofs are deferred to Appendices B and C.
A-B1 Case I: With Lipschitz Continuity Only for Loss Function
We first inspect the case with Assumptions 7, 9, 10, 11 and 12 imposed, which allows the loss function to be non-convex. By defining
and
$p^{t}=\left\{\begin{aligned} &{0,}&&{\eta^{t}\leqslant\frac{1}{L}}\\ &{1,}&&{\eta^{t}>\frac{1}{L}}\end{aligned}\right$., (40)
the following results can be expected.
Theorem 3
With $C_{\alpha}>0.5$, there is
$\displaystyle\quad\sum_{t=1}^{T}\frac{\eta^{t}}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
$\displaystyle\leqslant\frac{2\mathbb{E}\left\{F(\bm{w}^{1})-F(\bm{w}^{T})\right\}}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}+\frac{1}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}\sum_{t=1}^{T}\eta^{t}\Delta^{t}$
$\displaystyle\quad+\sum_{t=1}^{T}\frac{2(\eta^{t}+p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})\epsilon^{2}}{(2C_{\alpha}-1)^{2}\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}$
$\displaystyle\quad+\sum_{t=1}^{T}\frac{8(C_{\alpha})^{2}(\eta^{t}\sigma^{2}+(p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})(G^{2}+\sigma^{2}))}{(2C_{\alpha}-1)^{2}\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}},$ (41)
where
$\displaystyle\Delta^{t}=$ $\displaystyle\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\Bigg(\frac{2\alpha_{m}L^{2}(G^{2}+\sigma^{2})}{K^{t}}\sum_{k=2}^{K^{t}}(k-1)$
$\displaystyle\quad\cdot\sum_{i=1}^{k-1}(\eta_{m}^{t,i})^{2}+2\alpha_{m}\theta^{2}\Bigg).$ (42)
Proof:
Please refer to Appendix B. ∎
A-B2 Case II: With Lipschitz Continuity and Strong Convexity for Loss Function
For the case with Assumptions 7, 8, 9, 10, 11 and 12 imposed, which requires the loss function to be not only Lipschitz but also strongly-convex, the following results can be anticipated.
Theorem 4
With $0<\lambda^{t}<1$ and $C_{\alpha}>0.5$, the optimality gap $\mathbb{E}\left\{F(\bm{w}^{T})-F(\bm{w}^{*})\right\}$ is upper bounded in (4) as follows,
$\displaystyle\quad\mathbb{E}\left\{F(\bm{w}^{T})-F(\bm{w}^{*})\right\}$
$\displaystyle\leqslant\frac{L}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}\prod_{t=1}^{T-1}\gamma^{t}+\frac{L}{2}\sum_{t=1}^{T-1}\frac{(\eta^{t})^{2}}{\lambda^{t}}\Bigg(\Delta^{t}+$
$\displaystyle\quad\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}\Bigg)\cdot\prod_{i=t}^{T-1}(\gamma^{i+1})^{q^{i+1}},$ (43)
where
$q^{t}=\left\{\begin{aligned} &{0,}&&{t=T}\\ &{1,}&&{t<T}\end{aligned}\right$.,
and
Proof:
Please refer to Appendix C. ∎
Appendix B Proof of Theorem 3
@@
With Assumptions 7, 9, 10, 11, 12, and $C_{\alpha}>0.5$, the term $\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$ can be upper bounded as
$\displaystyle\quad\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}\leqslant\Delta^{t}+\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (44)
@@
Lemma 2
With Assumptions 9, 10, 12, and $C_{\alpha}>0.5$, the term $\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}$ can be bounded as
$\displaystyle 0\leqslant\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (45)
Proof:
Please refer to Appendix E. ∎
@@
$\displaystyle=\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}-\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
$\displaystyle\quad+\frac{(\eta^{t})^{2}L-\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}$ (46c)
$\displaystyle=\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}-\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
$\displaystyle\quad+\frac{(\eta^{t})^{2}L-\eta^{t}}{2}p^{t}\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}$ (46d)
$\displaystyle\leqslant-\frac{\eta^{t}}{2}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{\eta^{t}}{2}\Delta^{t}$
$\displaystyle\quad+\frac{(\eta^{t}+p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})\epsilon^{2}}{(2C_{\alpha}-1)^{2}}$
$\displaystyle\quad+\frac{4(C_{\alpha})^{2}(\eta^{t}\sigma^{2}+(p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})(G^{2}+\sigma^{2}))}{(2C_{\alpha}-1)^{2}}.$ (46e)
@@
$\displaystyle\quad+\sum_{t=1}^{T}\frac{2(\eta^{t}+p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})\epsilon^{2}}{(2C_{\alpha}-1)^{2}\sum_{t^{\prime}=1}^{T}\eta^{t}}$
$\displaystyle\quad+\sum_{t=1}^{T}\frac{8(C_{\alpha})^{2}(\eta^{t}\sigma^{2}+(p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})(G^{2}+\sigma^{2}))}{(2C_{\alpha}-1)^{2}\sum_{t^{\prime}=1}^{T}\eta^{t}}.$ (47)
@@
Appendix C Proof of Theorem 4
@@
$\displaystyle=\frac{L}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{t+1}-\bm{w}^{*}\right\rVert^{2}\right\}.$ (48b)
For the term $\mathbb{E}\left\{\left\lVert\bm{w}^{t+1}-\bm{w}^{*}\right\rVert^{2}\right\}$, with $0<\lambda^{t}<1$, it leads to
@@
$\displaystyle=\mathbb{E}\Bigl\{\big\lVert\bm{w}^{t+1}-\bm{w}^{t}+\eta^{t}\nabla F(\bm{w}^{t})$
$\displaystyle\qquad\qquad\qquad\qquad+\bm{w}^{t}-\eta^{t}\nabla F(\bm{w}^{t})-\bm{w}^{*}\big\rVert^{2}\Bigr\}$ (49a)
$\displaystyle\leqslant\frac{1}{\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t+1}-\bm{w}^{t}+\eta^{t}\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
$\displaystyle\quad+\frac{1}{1-\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t}-\eta^{t}\nabla F(\bm{w}^{t})-\bm{w}^{*}\right\rVert^{2}\right\}$ (49b)
$\displaystyle=\frac{(\eta^{t})^{2}}{\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{1}{1-\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t}-\bm{w}^{*}\right\rVert^{2}\right\}$
$\displaystyle\quad-\frac{2\eta^{t}}{1-\lambda^{t}}\left\langle\bm{w}^{t}-\bm{w}^{*},\nabla F(\bm{w}^{t})-\nabla F(\bm{w}^{*})\right\rangle$
@@
$\displaystyle\leqslant\frac{(\eta^{t})^{2}}{\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
$\displaystyle\quad+\frac{1-2\mu\eta^{t}+L^{2}(\eta^{t})^{2}}{1-\lambda^{t}}\mathbb{E}\left\{\left\lVert\bm{w}^{t}-\bm{w}^{*}\right\rVert^{2}\right\},$ (49d)
@@
  • the inequality (49d) is bounded by Assumptions 7 and 8.
@@
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{w}^{T}-\bm{w}^{*}\right\rVert^{2}\right\}\leqslant\prod_{t=1}^{T-1}\gamma^{t}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}$
$\displaystyle\quad+\sum_{t=1}^{T-1}\frac{(\eta^{t})^{2}}{\lambda^{t}}\prod_{i=t}^{T-1}(\gamma^{i+1})^{q^{i+1}}\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}.$ (50)
With the help of Lemma 1, and combining (48) and (C), there is
@@
$\displaystyle\leqslant\frac{L}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}\prod_{t=1}^{T-1}\gamma^{t}+\frac{L}{2}\sum_{t=1}^{T-1}\frac{(\eta^{t})^{2}}{\lambda^{t}}\Bigg[\Delta^{t}+$
$\displaystyle\quad\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}\Bigg]\cdot\prod_{i=t}^{T-1}(\gamma^{i+1})^{q^{i+1}}.$ (51)
@@
Appendix D Proof of Lemma 1